# https://github.com/NVlabs/OmniDrive

### v1

@@ 6 Conclusion
We address the challenges of end-to-end autonomous driving with LLM-agents by proposing OmniDrive-Agent and OmniDrive-nuScenes. OmniDrive-Agent adopts a novel Q-Former3D MLLM architecture that can efficiently handle high resolution multi-view videos. Our model design enables minimal adjustments to leverage 2D pre-trained knowledge while gaining important 3D spatial understanding. We additionally provide a novel benchmark for end-to-end autonomous driving which features counterfactual reasoning alongside 3D spatial awareness and reasoning tasks. OmniDrive-Agent demonstrates the efficacy by addressing high-resolution multi-view video input and illustrate excellent scene description and counterfactual reasoning. The model also yields compelling performance on open-loop 3D planning.
Limitations. Our method has not been validated on even larger datasets e.g., nuPlan [5]. The simulation of counterfactual outcomes, despite moving beyond single trajectories, does not yet consider reaction from other agents. This part can be further formed as a closed-loop setup, and we will leave it for future work.

### v2

@@ 6 Conclusion
We present OmniDrive, a holistic framework designed to advance end-to-end autonomous driving using LLM-agents. By introducing a counterfactual-based 3D driving Q&A pipeline, we enable scalable, high-quality data generation that significantly enhances decision-making capabilities. Pre-trained models on OmniDrive exhibit significant improvements on the DriveLM QA benchmark and nuScenes open-loop planning, underscoring the effectiveness and quality of our dataset. Furthermore, our exploration of two advanced frameworks, Omni-L and Omni-Q, provides valuable insights into the design of effective LLM-agents, highlighting the advantages of vision-language alignment in 3D spaces. These frameworks demonstrate the potential for improved reasoning and perception by integrating language models with 3D environmental understanding.
Limitations. The simulation of counterfactual outcomes, despite moving beyond single trajectories, does not yet consider reactions from other agents. As research on closed-loop planning simulators progresses, we aim to use closed-loop results to enhance effectiveness.
7 Acknowledgments
The team would like to give special thanks to the NVIDIA TSE Team, including Le An, Chengzhe Xu, Yuchao Jin, and Josh Park, for their exceptional work on the TensorRT deployment of OmniDrive.